A Bibliography of Publications of Tim Hopkins

Tim Hopkins

University of Kent at Canterbury

CT2 7NF U.K.

Tel: (+44) 1227-475430 FAX: (+44) 1227-762811

E-mail: trh@ukc.ac.uk (Internet)

13 October 2017 Version 1.09

Abstract

This bibliography records publications of Tim Hopkins.

Title word cross-reference

AXB^{T} + CXD^{T} = E [54].

2-Point [4].

539 [56].

705 [54]. 77 [54].

830 [56].

Acceleration [53, 58]. ACM [55, 43]. ADI [12]. Algebra [48, 47]. Algorithm [13, 15, 32, 45, 55, 56, 54, 63, 64]. Algorithms [35, 36, 22, 27, 30, 43, 44]. ANSI [16]. Applied [53, 58, 44, 63, 64]. Approximants [11]. approximate [55].

Arrays [29, 28]. AS030 [64]. AS057 [63]. AS135 [13]. AS188 [14]. AS193 [15]. ASR49 [13]. assignment [55].

B [6]. B-spline [6]. BASIC [8, 16]. Basis [6]. Better [18]. Boundary [4].

C [55]. Calculations [29, 28]. Case [65]. Christoffel [5]. Clusters [48, 47, 51]. Coefficients [11]. Collocation [3]. Comment [52, 61]. Comparative [33, 34]. Comparison [53, 58, 1, 2, 3]. Computer [66]. computers [62]. Configurable [21, 25, 26, 42]. Conjugate [49, 50]. Conjugate-Gradient [49]. Conjugate-Gradients [50]. Corrections [10]. coverage [57]. Cumulative [43, 44].

Data [23, 24]. Data-flow [23, 24]. defining [11]. Dependence [14]. Determining [5]. Developments [66]. Difference [4]. Differential [38, 39, 1, 2, 3]. Dimensional [12, 25, 26]. Distributing [28, 29].

Economical [4]. Eigenvalue [29, 28].

Eispack [61, 52]. Elliptic [12, 22, 25, 26, 27]. environments [51]. Epsilon [52, 61]. equation [54]. Equations [12, 33, 34, 35, 36, 37, 38, 39, 40, 41, 46, 60, 22, 25, 26, 27, 1, 2, 3, 10, 23, 24, 59, 62]. Estimates [13]. Estimation [14]. Evolution [65]. Exploiting [18].

Family [4]. far [31]. Finite [4]. Fitting [31]. flow [23, 24]. Fortran [55, 54, 60, 10]. Fortran-77 [54]. Functions [6].

G [55]. Galerkin [3, 6]. Generally [21, 25, 26, 42]. Givens [56]. GMRES [46, 59]. Gradient [49]. Gradients [50]. Graphical [20]. Graphics [17, 18, 31]. GRASP [55]. Guide [60, 19].

Half [64]. Half-Normal [64]. Hopscotch [22, 12, 27].

Implementation [17, 46, 21, 25, 26, 42, 59]. Improving [57]. Index [30, 43, 44]. Interface [20]. Interpolation [9]. Iterative [33, 34, 40, 41, 46, 60, 62, 59].

Kufarev [5].

L [55]. LAPACK [57]. Library [19]. Linear [33, 34, 35, 36, 37, 40, 41, 46, 48, 60, 22, 27, 13, 10, 23, 24, 47, 59, 4, 62]. Lines [3].

M [55]. Machine [52, 61]. Math [55]. Matrix [29, 28, 54]. Max [13]. Method [46, 53, 58, 3, 5, 59]. Methods [12, 33, 34, 40, 41, 49, 50, 60, 1, 2, 4, 6, 19, 62]. Microprocessors [7]. Min [13]. Min-Max [13]. Minimal [16]. Mixed [12]. modified [56]. MR [55]. Multidimensional [63]. Multigrid [21, 25, 26, 42]. Multiple [13].

NAG [19, 31]. Nets [34]. Network [25, 26]. Networks [33, 38, 39, 40, 41, 50, 21, 42]. no [55]. Non [4]. Non-linear [4].

Nonsymmetric [46, 59]. Normal [64]. Note [4]. Numerical [7, 1, 2, 19].

Order [4, 14]. Overrelaxation [35, 36].

P [55]. package [60, 62, 54]. Pade [11]. Parabolic [12]. Parallel [35, 36, 37, 38, 39, 40, 41, 46, 49, 50, 60, 22, 27, 62, 51, 59]. Parameters [5]. Pardalos [55]. Partial [38, 39, 1, 2, 3]. PBASIC [8, 16]. Peripheral [12]. PIM [60, 62]. Pitsoulis [55]. Plotting [64]. Point [4]. Portable [7]. Porting [47, 48]. Power [11]. Practice [19]. Preconditioned [49, 50]. Printing [63]. Problems [13, 4, 55]. Program [16]. programming [51]. Provide [18].

quadratic [55]. Quadrature [6]. Quasilinear [1, 2].

Rational [9]. Regression [13]. Reliable [9]. Remark [13, 32, 45, 63, 64, 54, 55, 56]. Resende [55]. respect [11]. Restarted [46, 59]. Revised [15]. Routine [52, 61]. Routines [31]. Rules [6].

S [55]. Schwarz [5]. Schwarz-Christoffel [5]. Science [66]. Sensitivity [11]. Sequences [14]. Series [11]. Sets [10]. Sixth [4]. Software [18, 7, 65, 55, 54]. Solution [37, 38, 39, 40, 41, 22, 25, 26, 27, 1, 2, 10, 24, 62, 55]. Solutions [23]. Solve [33, 34]. solving [54]. Some [6, 34]. SOR [53, 58]. sparse [55]. Spectral [15]. Spline [31, 6]. standard [56]. Statistical [30]. Statistics [44, 63, 64]. Story [31]. Study [33, 34, 65]. Subroutines [48, 10, 47, 55]. Sylvester [54]. Systems [33, 34, 35, 36, 37, 40, 41, 46, 60, 23, 24, 59, 62].

Tables [63]. Teaching [66]. Techniques [53, 58]. Term [12]. Test [15, 57]. their [11]. Three [25, 26]. Three-Dimensional [25, 26]. Toeplitz [10]. Toolpack [17, 20].

Toolpack/1 [17, 20]. Tools [18]. Trans [55]. transformations [56]. Transputer [33, 34, 38, 39, 40, 41, 50, 21, 25, 26, 42, 23, 24, 29, 28]. Transputers [48, 47]. Triangular [37]. Two [12].

Use [17]. User [60]. Using [51, 6, 33, 34, 40, 41, 22, 27, 55].

Value [4]. Verifier [8, 16]. version [60]. visit [56].

Workstation [17, 18, 31]. Workstations [48, 47, 51].

References

Hopkins:1976:CNM

[1] T. R. Hopkins and R. Wait. A comparison of numerical methods for the solution of quasilinear partial differential equations. Comput. Methods Appl. Mech. Engrg., 9: 181–190, 1976.

Hopkins:1977:CNM

[2] T. R. Hopkins. A Comparison of Numerical Methods for the Solution of Quasilinear Partial Differential Equations. PhD thesis, University of Liverpool, Liverpool, U.K., February 1977.

Hopkins:1978:CGC

[3] T. R. Hopkins and R. Wait. A comparison of Galerkin, collocation and the method of lines for partial differential equations. Internat. J. Numer. Methods Engrg., 12: 1081–1107, 1978.

Hopkins:1978:NFE

[4] T. R. Hopkins. Note on a family of economical sixth order finite difference methods for non-linear 2-point boundary value problems. BIT, 18:497–499, 1978.

Hopkins:1979:KMD

[5] T. R. Hopkins and D. E. Roberts. Kufarev’s method for determining the Schwarz-Christoffel parameters. Numer. Math., 33:353–365, 1979.

Hopkins:1979:SQR

[6] T. R. Hopkins and R. Wait. Some quadrature rules for Galerkin methods using Bspline basis functions. Comput. Methods Appl. Mech. Engrg., 19:401–416, 1979.

Genz:1980:PNS

[7] A. C. Genz and T. R. Hopkins. Portable numerical software for microprocessors. In M. A. Hennell and L. M. Delves, editors, Production and Assessment of Numerical Software. Academic Press, New York, 1980.

Hopkins:1980:PVB

[8] T. R. Hopkins. PBASIC – A verifier for BASIC. Software — Practice and Experience, 10:175–182, 1980.

Graves-Morris:1981:RRI

[9] P. R. Graves-Morris and T. R. Hopkins. Reliable rational interpolation. Numer. Math., 36:111–128, 1981.

Hopkins:1981:CFS

[10] T. R. Hopkins. Corrections to Fortran subroutines for the solution of Toeplitz sets of linear equations. IEEE Trans. Acoust. Speech Signal Process., 29:1212, 1981.

Hopkins:1982:SCP

[11] T. R. Hopkins. On the sensitivity of the coefficients of Pade approximants with respect to their defining power series coefficients. J. Comput. Appl. Math., 8:105– 109, 1982.

Broom:1983:PAH

[12] M. A. Broom and T. R. Hopkins. Peripheral and ADI Hopscotch methods for two dimensional parabolic and elliptic equations with a mixed term. J. Comput. Appl. Math., 9:239–246, 1983.

Grant:1983:ARA

[13] P. M. Grant and T. R. Hopkins. ASR49 – A remark on algorithm AS135 : Minmax estimates for linear multiple regression problems. J. Roy. Statist. Soc. Ser. C, 32:345–347, 1983.

Hopkins:1983:AEO

[14] T. R. Hopkins, P. H. Welch, and J. Kollerstrom. AS188 – estimation of the order of dependence in sequences. J. Roy. Statist. Soc. Ser. C, 32:185–196, 1983.

Hopkins:1983:ARA

[15] T. R. Hopkins. AS193 – A revised algorithm for the spectral test. J. Roy. Statist. Soc. Ser. C, 32:328–335, 1983.

Hopkins:1983:PVP

[16] T. R. Hopkins. PBASIC – A verifier program for ANSI minimal Basic. ACM Trans. Math. Software, 9:391–394 (plus microfiche listing), 1983.

Broom:1987:IUT

[17] M. A. Broom and T. R. Hopkins. The implementation and use of Toolpack/1 on a graphics workstation. Software — Practice and Experience, 17:561–573, 1987.

Broom:1988:EWG

[18] M. A. Broom, S. P. Holmes, and T. R.

Hopkins. Exploiting workstation graphics to provide better software tools. In Proceedings of the International Workshop on Software Engineering and Its Applications Volume 2, Toulouse, pages 219–230. Ec2, 1988. ISBN 2-906899-16-X.

Hopkins:1988:NMP

[19] T. R. Hopkins and C. Phillips. Numerical Methods in Practice – A Guide to the NAG Library. Addison-Wesley, Reading, MA, USA, 1988.

Broom:1989:GIT

[20] M. A. Broom, S. P. Holmes, and T. R. Hopkins. A graphical interface for Toolpack/1. Information and Software Technology, 31:305–312, 1989.

El-Giar:1989:GCMa

[21] O. El-Giar and T. R. Hopkins. A generally configurable multigrid implementation for transputer networks. In J. Wexler, editor, Developing Transputer Applications, pages 80–88. IOS, Amsterdam, 1989. Available as a compressed Postscript file from unix.hensa.

ac.uk in the file pub/misc/ukc.
reports/comp.sci/reports/23-92.ps.

Z; abstract in the file pub/misc/ukc.

reports/comp.sci/abstracts/23-92.

El-Giar:1989:PSL

[22] Mohamed Osama El-Giar and Tim Hopkins. The parallel solution of linear elliptic equations using hopscotch algorithms. Technical Report 62, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., June 1989. Published as [27]. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/62.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/62.

Hopkins:1989:TDFa

[23] T. R. Hopkins and P. H. Welch. Transputer data-flow solutions for systems of linear equations. In G. R. Joubert, D. J. Evans, and F. J. Peters, editors, Parallel Computing ’89 – Proceedings of the International Conference, Leiden, pages 195–200. North Holland-Elsevier Science Publishers, Amsterdam, New York, Oxford, 1989. An extended version of this paper is available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/68.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/68.

Hopkins:1989:TDFb

[24] Tim Hopkins and Peter Welch. Transputer data-flow solution for systems of linear equations. Technical Report 68, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., September 1989. An extended version of [23]. Available as a compressed Postscript file

from unix.hensa.ac.uk in the file pub/
misc/ukc.reports/comp.sci/reports/

68.ps.Z; abstract in the file pub/misc/

ukc.reports/comp.sci/abstracts/68.

El-Giar:1990:GCMa

[25] O. El-Giar and T. R. Hopkins. A generally configurable multigrid implementation for the solution of three-dimensional elliptic equations on a transputer network. In T. L. Kunii and D. May, editors, Transputer/occam Japan 3, pages 35–45. IOS, Amsterdam, 1990. Available as a compressed Postscript file

from unix.hensa.ac.uk in the file pub/
misc/ukc.reports/comp.sci/reports/

79.ps.Z; abstract in the file pub/misc/

                                        ukc.reports/comp.sci/abstracts/79.

El-Giar:1990:GCMb

[26] Mohamed Osama El-Giar and Tim Hopkins. A generally configurable multigrid implementation for the solution of three-dimensional elliptic equations on a transputer network. Technical Report 79, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., October 1990. Published as [25]. Available as a compressed Postscript file from unix.hensa.ac.uk

                                        in the file pub/misc/ukc.reports/comp.
                                        sci/reports/79.ps.Z; abstract in the
                                        file pub/misc/ukc.reports/comp.sci/
                                        abstracts/79.

El-Giar:1990:PSL

[27] O. El-Giar and T. R. Hopkins. The parallel solution of linear elliptic equations using Hopscotch algorithms. In L. Freeman and C. Phillips, editors, Applications of Transputers 1 – Proceedings of the First International Conference on Applications of Transputers, Liverpool, pages 151–159. IOS, Amsterdam, 1990. Available as a compressed Postscript file from unix.hensa.ac.uk

                                        in the file pub/misc/ukc.reports/comp.
                                        sci/reports/62.ps.Z; abstract in the
                                        file pub/misc/ukc.reports/comp.sci/
                                        abstracts/62.

Hopkins:1990:DMEa

[28] T. R. Hopkins and B. J. Vowden. Distributing matrix eigenvalue calculations over transputer arrays. In H. S. M. Zedan, editor, Real-time Systems with Transputers, pages 305–312. IOS, Amsterdam, 1990. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/81.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/81.

Hopkins:1990:DMEb

[29] Tim Hopkins and Barry Vowden. Distributing matrix eigenvalue calculations over transputer arrays. Technical Report 81, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., October 1990. Published as [28]. Available as a compressed Postscript file

from unix.hensa.ac.uk in the file pub/
misc/ukc.reports/comp.sci/reports/

81.ps.Z; abstract in the file pub/misc/

ukc.reports/comp.sci/abstracts/81.

Hopkins:1990:ISA

[30] T. R. Hopkins and D. R. Morse. Index of statistical algorithms. J. Roy. Statist. Soc. Ser. C, 39(1):177–187, 1990.

Hopkins:1990:NSF

[31] Tim Hopkins. NAG spline fitting routines on a graphics workstation – the story so far. Technical Report 76, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., July 1990. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/76.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/76.

Hopkins:1990:RA

[32] T. R. Hopkins and D. R. Morse. Remark on algorithm 620. ACM Trans. Math. Software, 16(4):401–403, December 1990.

Cunha:1991:CSI

[33] Rudnei Dias da Cunha and Tim Hopkins. A comparative study of iterative methods to solve systems of linear equations using transputer networks. Technical Report 89, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., June 1991. Published as [34]. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/89.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/89.

Cunha:1991:CSS

[34] R. D. da Cunha and T. R. Hopkins. A comparative study of some iterative methods to solve systems of linear equations using transputer nets. In J. Edwards, editor, Occam and the Transputer – Current Developments, pages 85–99. IOS, Amsterdam, 1991. Available as a compressed Postscript file

from unix.hensa.ac.uk in the file pub/
misc/ukc.reports/comp.sci/reports/

89.ps.Z; abstract in the file pub/misc/

ukc.reports/comp.sci/abstracts/89.

Cunha:1991:POAa

[35] R. D. da Cunha and T. R. Hopkins. Parallel overrelaxation algorithms for systems of linear equations. In T. L. Kunii P. Welch, D. Stiles and A. Bakkers, editors, Transputing ’91, pages 159–169. IOS, Amsterdam, 1991. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/80.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/80.

Cunha:1991:POAb

[36] Rudnei Dias da Cunha and Tim Hopkins.

Parallel overrelaxation algorithms for systems of linear equations. Technical Report 80, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., January 1991. Published as [35]. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/80.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/80.

Cunha:1991:PST

[37] Rudnei Dias da Cunha and Tim Hopkins. The parallel solution of triangular systems of linear equations. Technical Report 86, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., June 1991. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/86.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/86.

Cunha:1992:PSPa

[38] R. D. da Cunha and T. R. Hopkins.

The parallel solution of partial differential equations on transputer networks. In J. Luo and G. L. Reijns, editors, Transputing for Numerical and Neural Network Applications, pages 96–109. IOS, Amsterdam, 1992. Available as a compressed Postscript file from unix.hensa.

ac.uk in the file pub/misc/ukc.
reports/comp.sci/reports/17-92.ps.

Z; abstract in the file pub/misc/ukc.

reports/comp.sci/abstracts/17-92.

Cunha:1992:PSPb

[39] Rudnei Dias da Cunha and Tim Hopkins. The parallel solution of partial differential equations on transputer networks. Technical Report 17-92, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., June 1992. Published as [38]. Available as a compressed Postscript

file from unix.hensa.ac.uk in the file
pub/misc/ukc.   reports/comp.sci/
reports/17-92.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/17-92.

Cunha:1992:PSSa

[40] R. D. da Cunha and T. R. Hopkins. The parallel solution of systems of linear equations using iterative methods on transputer networks. In J. Luo and G. L. Reijns, editors, Transputing for Numerical and Neural Network Applications, pages 1–13. IOS, Amsterdam, 1992. Available as a compressed Postscript

file from unix.hensa.ac.uk in the file
pub/misc/ukc.   reports/comp.sci/
reports/16-92.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/16-92.

Cunha:1992:PSSb

[41] Rudnei Dias da Cunha and Tim Hopkins. The parallel solution of systems of linear equations using iterative methods on transputer networks. Technical Report 16-92, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., June 1992. Published as [40]. Available as a compressed Postscript

file from unix.hensa.ac.uk in the file
pub/misc/ukc.   reports/comp.sci/
reports/16-92.ps.Z; abstract in the

file pub/misc/ukc.reports/comp.sci/
abstracts/16-92.

El-Giar:1992:GCM

[42] Osama El-Giar and Tim Hopkins. A generally configurable multigrid implementation for transputer networks. Technical Report 23-92, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., September 1992. Published as [21]. Available as a compressed Postscript

file from unix.hensa.ac.uk in the file
pub/misc/ukc.   reports/comp.sci/
reports/23-92.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/23-92.

Hopkins:1992:CIAa

[43] Tim Hopkins and David Morse. Cumulative index to the ACM algorithms. Technical Report 24-92, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., October 1992. Available as a compressed Postscript file from unix.hensa.

ac.uk in the file pub/misc/ukc.
reports/comp.sci/reports/24-92.ps.

Z; abstract in the file pub/misc/ukc.

reports/comp.sci/abstracts/24-92.

Hopkins:1992:CIAb

[44] Tim Hopkins and David Morse. Cumulative index to the applied statistics algorithms. Technical Report 25-92, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., October 1992. Available as a compressed Postscript file from unix.hensa.

ac.uk in the file pub/misc/ukc.
reports/comp.sci/reports/25-92.ps.

Z; abstract in the file pub/misc/ukc.

reports/comp.sci/abstracts/25-92.

Hopkins:1992:RA

[45] T. R. Hopkins. Remark on algorithm 540. ACM Trans. Math. Software, 18(3):343– 344, September 1992.

Cunha:1993:PIR

[46] Rudnei Dias da Cunha and Tim Hopkins. A parallel implementation of the restarted GMRES iterative method for nonsymmetric systems of linear equations. Technical Report 7-93, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., May 1993. Published as [59]. Available as a compressed Postscript file from unix.hensa.ac.uk

                                        in the file pub/misc/ukc.reports/comp.
                                        sci/reports/7-93.ps.Z; abstract in the
                                        file pub/misc/ukc.reports/comp.sci/
                                        abstracts/7-93.

Cunha:1993:PLAa

[47] R. D. da Cunha and T. R. Hopkins. Porting linear algebra subroutines from transputers to clusters of workstations. In R. Grebe, J. Hektor, S. C. Hilton, M. R. Jane, and P. H. Welch, editors, Proceedings of the World Transputer Congress, pages 660–667. IOS, Amsterdam, 1993. Available as a compressed Postscript file from unix.hensa.ac.uk

                                        in the file pub/misc/ukc.reports/comp.
                                        sci/reports/6-93.ps.Z; abstract in the
                                        file pub/misc/ukc.reports/comp.sci/
                                        abstracts/6-93.

Cunha:1993:PLAb

[48] Rudnei Dias da Cunha and Tim Hopkins. Porting linear algebra subroutines from transputers to clusters of workstations. Technical Report 6-93, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., May 1993. Published as [47]. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/6-93.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/6-93.

Cunha:1993:PPCa

[49] R. D. da Cunha and T. R. Hopkins. Parallel preconditioned conjugategradient methods. Transputer Communications, 1(2):111–125, November 1993. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.   reports/
comp.sci/reports/5-93.ps.Z; abstract
in the file pub/misc/ukc.reports/comp.
sci/abstracts/5-93.

Cunha:1993:PPCb

[50] Rudnei Dias da Cunha and Tim Hopkins. Parallel preconditioned conjugategradients methods on transputer networks. Technical Report 5-93, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., April 1993. Published as [49]. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/5-93.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/5-93.

Cunha:1993:UPP

[51] R. D. da Cunha and Tim Hopkins. Using parallel programming environments on clusters of workstations. Technical Report 12-93, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., December 1993. Available as a compressed Postscript file from unix.hensa.

                                        ac.uk in the file pub/misc/ukc.
                                        reports/comp.sci/reports/12-93.ps.

Z; abstract in the file pub/misc/ukc.

                                        reports/comp.sci/abstracts/12-93.

Hopkins:1993:CEM

[52] T. R. Hopkins and J. B. Slater. A comment on the Eispack machine epsilon routine. ACM SIGNUM, 28(4):2– 6, October 1993. Available as a compressed Postscript file from unix.hensa.

                                        ac.uk in the file pub/misc/ukc.
                                        reports/comp.sci/reports/18-94.ps.

Z; abstract in the file pub/misc/ukc.

                                        reports/comp.sci/abstracts/18-94.

Cunha:1994:CATa

[53] R. D. da Cunha and T. R. Hopkins. A comparison of acceleration techniques applied to the SOR method. In A. Cuyt, editor, Proceedings on Nonlinear Numerical Methods and Rational Approximation, pages 247–260. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, March 1994. Published as [53]. Available as a compressed Postscript file from unix.hensa.ac.uk

                                        in the file pub/misc/ukc.reports/comp.
                                        sci/reports/4-94.ps.Z; abstract in the
                                        file pub/misc/ukc.reports/comp.sci/
                                        abstracts/4-94.

Hopkins:2002:RAF

[54] Tim Hopkins. Remark on Algorithm 705: A Fortran-77 software package for solving the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Transactions on Mathematical Software, 28(3): 372–375, 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hopkins:2003:RAF

[55] Tim Hopkins. Remark on: “Algorithm 769: Fortran subroutines for approximate solution of sparse quadratic assignment problems using GRASP” [ACM Trans. Math. Software 23 (1997), no. 2, 196– 208; MR 1671740] by P. M. Pardalos, L. S. Pitsoulis, and M. G. C. Resende. ACM Transactions on Mathematical Software, 29(3):349–351, 2003. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hanson:2004:AAV

[56] Richard J. Hanson and Tim Hopkins. Algorithm 830: another visit with standard and modified Givens transformations and a remark on Algorithm 539. ACM Transactions on Mathematical Software, 30(1): 86–94, 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Barnes:2007:ITC

[57] David J. Barnes and Tim R. Hopkins. Improving test coverage of LAPACK. Applicable algebra in engineering, communication and computing, 18(3):209–222, 2007. CODEN AAECEW. ISSN 0938-1279 (print), 1432-0622 (electronic).

Cunha:1994:CATb

[58] R. D. da Cunha and Tim Hopkins.

A comparison of acceleration techniques applied to the SOR method. Technical Report 4-94, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., March 1994. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/4-94.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/4-94.

Cunha:1994:PIR

[59] R. D. da Cunha and T. R. Hopkins. A parallel implementation of the restarted GMRES iterative method for nonsymmetric systems of linear equations. Advances in Computational Mathematics, 2(3):261–277, 1994. Available as a compressed Postscript file from unix.

hensa.ac.uk in the file pub/misc/ukc.
reports/comp.sci/reports/7-93.ps.

Z; abstract in the file pub/misc/ukc.

reports/comp.sci/abstracts/7-93.

Cunha:1994:PPI

[60] R. D. da Cunha and Tim Hopkins. PIM 1.0: The parallel iterative methods package for systems of linear equations. User’s guide (Fortran 77 version). Technical Report 2-94, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., January 1994. Available as a compressed Postscript file from unix.hensa.ac.uk

in the file pub/misc/ukc.reports/comp.
sci/reports/2-94.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/2-94.

Hopkins:1994:CEM

[61] Tim Hopkins and John Slater. A comment on the eispack machine epsilon routine. Technical Report 18-94, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., September 1994. Published as [52]. Available as a compressed Postscript

file from unix.hensa.ac.uk in the file
pub/misc/ukc.   reports/comp.sci/
reports/18-94.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/18-94.

Hopkins:1994:PIM

[62] Tim Hopkins and Rudnei Dias da Cunha.

The parallel iterative methods (PIM) package for the solution of systems of linear equations on parallel computers. Technical Report 19-94, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., September 1994. Available as a compressed Postscript file from unix.hensa.

ac.uk in the file pub/misc/ukc.
reports/comp.sci/reports/19-94.ps.

Z; abstract in the file pub/misc/ukc.

reports/comp.sci/abstracts/19-94.

Hopkins:1994:RASa

[63] Tim Hopkins and David Morse. Remark on Applied Statistics Algorithm AS057 printing multidimensional tables. Technical Report 29-94, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., December 1994. Available as a compressed Postscript file from unix.hensa.

ac.uk in the file pub/misc/ukc.
reports/comp.sci/reports/29-94.ps.

Z; abstract in the file pub/misc/ukc.

reports/comp.sci/abstracts/29-94.

Hopkins:1994:RASb

[64] Tim Hopkins and David Morse. Remark on Applied Statistics Algorithm AS030 half-normal plotting. Technical Report 28-94, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., December 1994. Available as a compressed Postscript

file from unix.hensa.ac.uk in the file
pub/misc/ukc.   reports/comp.sci/
reports/28-94.ps.Z; abstract in the
file pub/misc/ukc.reports/comp.sci/
abstracts/28-94.

Hopkins:1994:SEC

[65] Tim Hopkins. Software evolution: A case study. Technical Report 27-94, Computing Laboratory, University of Kent at Canterbury, Canterbury, Kent CT2 7NF, U.K., December 1994. Available as a compressed Postscript file from unix.hensa.

                                        ac.uk in the file pub/misc/ukc.
                                        reports/comp.sci/reports/27-94.ps.

Z; abstract in the file pub/misc/ukc.

                                        reports/comp.sci/abstracts/27-94.

Bateman:1992:DTC

[66] D. C. Bateman and T. R. Hopkins, editors. Developments in the Teaching of Computer Science. Canterbury, Kent CT2 7NF, U.K., University of Kent at Canterbury, Canterbury, Kent, June 1992.